OpenStudy (anonymous):

Jerry is two less than three times his daughter's age. Jerry is 28; how old is his daughter? 4 6 8 10

OpenStudy (anonymous):

can i have a medal for this? Jerry = 28 (two less than 3 times his daughters age) so, (28 + 2) = 30, 30/3 = 10. Jerry's daughter is 10 years old.

OpenStudy (anonymous):

she is 10

OpenStudy (anonymous):

yep